OpenStudy (anonymous):

cos^2x = (1 - cos2x)/2 TRUE or FALSE

OpenStudy (anonymous):

False.

OpenStudy (zzr0ck3r):

so cos^2(x) = 1/2-cos^2(x)/2

OpenStudy (zzr0ck3r):

plug in pi for x and see if that is true 1 = 1/2 - 1/2 = 0

OpenStudy (zzr0ck3r):

@ayyebigkid please explain your answer, or guide them to an answer for them selves

OpenStudy (anonymous):

FALSE! oh thanks...

OpenStudy (anonymous):

Okay! Sorry, I am new. :P I will do my best to explain next time!

OpenStudy (anonymous):

But @zzr0ck3r what about this equation?

OpenStudy (anonymous):

cos((π/2) - x) = sinx I cannot plug in π again

OpenStudy (zzr0ck3r):

well plugging in only works to show they are not equal, it does not show that they are.

OpenStudy (anonymous):

Oh wait, never mind then! it is TRUE

OpenStudy (anonymous):

I see what you mean now :) Thanks again